OpenStudy (anonymous):

Solve for v, eliminating w: 7v - w = f av + bw = 2 the answer i got was (2+fb)/(a+7b)

OpenStudy (anonymous):

7v-w + w = f+w 7v=f+w v= f+w/7 That's what I got for the firsto ne.

OpenStudy (anonymous):

OH WAIT sorry.

OpenStudy (anonymous):

so am i wrong ?

OpenStudy (anonymous):

I don't know... I gave up when I had to multiply by b abd then divide by a and b and I don't know.

OpenStudy (anonymous):

ha okay thanks